OpenStudy (anonymous):

The local community park is designing a pool using the Golden Ratio. If the pool is 110 feet wide, about how long should it be?

OpenStudy (anonymous):

What is the golden ratio?

OpenStudy (pfenn1):

The golden ratio is also called the divine proportion. It equals 1.618 (approx).

OpenStudy (pfenn1):

Assuming the length is larger than the width 1.618=x/110

OpenStudy (anonymous):

does it have any thing to do with fibonacci numbers?

OpenStudy (anonymous):

^Yeah, that's how I would solve it too. :-)

OpenStudy (anonymous):

a/b = 1/2(1+sqrt(5))

OpenStudy (anonymous):

110 is not correct. :P

OpenStudy (pfenn1):

@drpeppergal , according to wikipedia, the ratio of sequential elements of the Fibonacci sequence approaches the golden ratio asymptotically.

OpenStudy (anonymous):

It's 177.98 @computergeek123 :-)

OpenStudy (anonymous):

1.618*110 = x

OpenStudy (anonymous):

Typo: it should be \[\LARGE \Phi^n = Fib(n) \Phi + Fib(n-1) \]

OpenStudy (anonymous):

Which side is which?

OpenStudy (anonymous):

68 (approx) 110/68 = 1.618

OpenStudy (anonymous):

so i woulld put 180?

OpenStudy (anonymous):

OK, assuming one is less than other, I see now...

OpenStudy (anonymous):

180 is too much rounding, I guess. 178 is better, but we are already rounding phi to 1.618 only :-)

OpenStudy (anonymous):

I still say there are 2 answers, why should you divide length by width?

OpenStudy (anonymous):

Sure, I agree @estudier. It's a pretty badly formulated question. @computergeek123 should pick between one of two. :-)

OpenStudy (anonymous):

well theres 100, 120, 150 and 180 so i think 180 is best? @bmp

OpenStudy (anonymous):

Yup, then it's

OpenStudy (anonymous):

OK, assumed in the choices......grrr.

OpenStudy (anonymous):

is this next one correct? Which number comes next in this sequence? 2, 5, 11, 23, 47 … Answer 68 82 95✔ 101

OpenStudy (anonymous):

Double +1

OpenStudy (anonymous):

Yeah, @estudier is correct. :-)

OpenStudy (anonymous):

The tick was a clue:-)

OpenStudy (anonymous):

Meg grows broccoli in a square plot in her garden. She created a drawing of her broccoli plot on a coordinate grid. The locations of three corners of the plot are represented by coordinates (3, 0), (1, 6), and (7, 8). Which of these coordinates represent the fourth corner? (8, 2) (8, 3) (9, 2)✔✔ (9, 3)

OpenStudy (anonymous):

Bibi is using a software program that allows her to change the shape of a two-dimensional figure by making changes to the sides. She starts with a quadrilateral with opposite sides parallel. She stretches the shorter edges of the quadrilateral to make all four sides congruent. Which statement best describes the figure that she created? Answer She changed a kite to a square. She changed a rectangle to a square. She changed a trapezoid to a rhombus.✔ She changed a parallelogram to a rhombus.

OpenStudy (anonymous):

The last one, it's just the definition of a trapezoid and a rhombus, so, it's correct :-)

OpenStudy (anonymous):

So i have the correct answers??